# Probabilistic Sampling of Balanced K-Means using Adiabatic Quantum Computing

### v1

@@ 3.5 Clustering as QUBO
To solve the clustering problem using AQC, a QUBO formulation is required. We use a variation of the one-hot encoding approach (Date et al., 2021; Arthur & Date, 2021) that uses a matrix ${Z}\in\{0,1\}^{K\times I}$ to encode the cluster assignment of $I$ samples to $K$ clusters. Each row corresponds to one of $K$ clusters and each column to one of $I$ samples. An entry ${Z}_{ki}=1$ indicates that the sample $x_{i}$ belongs to cluster $c_{k}$.
As each sample needs to be assigned to a single cluster, the sum of each column of $Z$ needs to satisfy the constraint $\sum_{k}{Z}_{ki}=1~{}\forall i$. The implementation of constrained clustering, where each cluster has a fixed size $s_{k}$ furthermore requires the row constraints $\sum_{i}{Z}_{ki}=s_{k}~{}\forall k$ on $Z$.
With the cost $q(i,j,k)$ for assigning the pair of samples $x_{i}$ and $x_{j}$ to the same cluster $c_{k}$, the optimization problem reads
$\hat{{Z}}=\operatorname*{arg\,min}_{Z}\sum_{k}\sum_{i}\sum_{j}Z_{ki}Z_{kj}q(i,j,k)$ (7)
[…]
@@
By vectorizing $Z$ in row-major order as $\mathbf{z}=\text{vec}(Z)$, the optimization problem can be rewritten in matrix form as
$\hat{\mathbf{z}}=\operatorname*{arg\,min}_{\mathbf{z}}\mathbf{z}^{\intercal}\,\mathbf{Q}\,\mathbf{z}\quad\text{s.t.}\quad\mathbf{G}\mathbf{z}=\mathbf{d},$ (8)
@@
where $\mathbf{Q}$ is a block diagonal matrix with blocks $\mathbf{Q}_{0},...,\mathbf{Q}_{K}$ and $\mathbf{G}\mathbf{z}=\mathbf{d}$ corresponds to the matrix formulation of the constraints. Each block $\mathbf{Q}_{k}$ of the cost matrix is a square matrix that contains the costs $q(i,j,k)$ at $\mathbf{Q}_{k,ij}$.
While the constraints are not directly covered by the QUBO problem, they can be modeled using Lagrangian multipliers. To avoid a mixed discrete and continuous optimization problem, a quadratic penalty reformulation $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}$ is chosen, leading to the minimization problem
$\hat{\mathbf{z}}=\operatorname*{arg\,min}_{\mathbf{z}}\mathbf{z}^{\intercal}\mathbf{Q^{\prime}}\mathbf{z}+\mathbf{b^{\prime}}^{\intercal}\mathbf{z}$ (9)
with
$\mathbf{Q^{\prime}}=\mathbf{Q}+\lambda\mathbf{G}^{\intercal}\mathbf{G}$ (10)
$\mathbf{b^{\prime}}=-2\lambda\mathbf{G}^{\intercal}\mathbf{b}.$ (11)
@@
In contrast to a linear penalty approach, where the multiplier $\lambda$ needs to be optimized, our selection only requires a sufficiently high $\lambda$, as all constraint violations result in an increased penalty term. With such selection, the penalty term evaluates to zero $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}=0$ if all constraints are fulfilled, and thus, the minimizer $\hat{\mathbf{z}}$ of the modified optimization problem is a minimizer of the original optimization problem in Equation equation 8.
Besides modeling balanced clustering, further constraints can be introduced using Lagrangian multipliers. This allows to represent a wide range of tasks as QUBO clustering problems solvable with AQC.
4 Probabilistic Quantum Clustering
4.1 Motivation
@@
Our clustering formulation employs a mixture of Gaussians to explain the observations, where the cluster assignments are latent variables. Each sample in a cluster $c_{k}$ is modeled as a sample drawn from a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{k},\mathbf{I})$ with mean $\mathbf{\mu}_{k}$ and identity covariance according to the k-means setting. Following a Bayesian approach, the best clustering solution can be found by maximizing the posterior probability over possible assignments $Z$ of points to clusters.
$p(Z|\mathbf{X})=\frac{p(\mathbf{X}|{Z})p({Z})}{\sum_{{Z}^{\prime}}p(\mathbf{X}|{Z^{\prime}})}=\frac{p(\mathbf{X}|{Z})p({Z})}{A}.$ (12)
@@
While this approach provides a probabilistic estimate by jointly modeling information about the possible cluster configurations and the distribution of data points, it is often intractable to evaluate due to the partitioning function $A=\sum_{{Z}^{\prime}}p(\mathbf{X}|{Z^{\prime}})$. Computing $A$ requires to sum over all possible solutions $Z$, which grows exponentially with the number of samples in the clustering problem. To overcome this, we utilize an AQC that samples directly from a Boltzmann distribution, which we parameterize according to the probabilistic clustering problem.
4.2 Data Model
@@
Determining the cost function of the optimization problem is a design choice of the algorithm. For our probabilistic approach, a well-defined data distribution, that forms the basis of the cost function is required. While many tasks approached in quantum computer vision, such as tracking (Zaech et al., 2022) or synchronization (Benkner et al., 2021; Birdal et al., 2021), costs are based on learned metrics or heuristics, they can also be trained to reflect the properties required in our approach.
@@
We therefore, follow the mixture of Gaussian model, where each cluster generates samples from a normal distribution. With the independence of observations and clusters, given the distribution parameters, the likelihood of the joint observations for an assignment $Z$ is given as the product of the individual likelihods
$f(\mathbf{X}|{Z})=\prod_{k=1}^{K}f(\mathbf{X}|{Z_{k}})=\prod_{k=1}^{K}\prod_{i\in Z_{k}}f(\mathbf{x}_{i}|{Z_{k}})=$ (13)
$\prod_{k=1}^{K}\prod_{i\in Z_{k}}\frac{1}{\sqrt{2\pi^{d}}}\mbox{exp}\left[-\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})\right]=$ (14)
$\frac{1}{\sqrt{2\pi^{d}}}\mbox{exp}\left[-\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})\right]$ (15)
@@
where the likelihood $f(\mathbf{x}_{i}|{Z_{k}})$ corresponds to a Gaussian distribution that follows $\mathcal{N}(\mathbf{\mu}_{k},\mathbf{I})$ and $d$ is the dimensionality of the space. This result can be used to formulate the energy-based model with the energy function
$E(\mathbf{X}|Z)=\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}})^{\intercal}\mathbf{I}({\mathbf{x}_{i}}-{\mathbf{\mu}_{k}}).$ (16)
@@
$p(Z|\mathbf{X})=\frac{\mbox{exp}\left[-(E(\mathbf{X}|Z)+E(Z))\right]}{\sum_{{Z}^{\prime}}\mbox{exp}\left[-E(\mathbf{X}|Z^{\prime})\right]}.$ (17)
@@
In this formulation, the energy of the assignment $E(Z)$ corresponds to the prior $p({Z})$, which models feasible and infeasible solutions by an indicator function. In the case of balanced clustering, this corresponds to allowing assignments that have each point assigned to exactly one cluster and a cluster size according to $s_{k}$.
Searching the most likely clustering solution corresponds to finding lowest energy solution on on the AQC, as well as to the Maximum-a-Posteriori $\hat{Z}_{\text{map}}$ estimate of the assignment
$\hat{Z}_{\text{map}}=\operatorname*{arg\,max}_{Z}p(Z|\mathbf{X})p({Z})=\operatorname*{arg\,min}_{Z}E(\mathbf{X}|Z)+E(Z).$ (18)
@@
As the AQC qubit system is an Ising model, it requires formulating $E(\mathbf{X}|Z)$ and $E(Z)$ quadratic in the optimization variables $Z$, enabling the joint discovery of assignments and cluster means. This is achieved by using the maximum likelihood estimator of the mean, resulting in a quadratic energy formulation that fits the Ising model in Equation 9
$E_{k}(\mathbf{X}|Z)=\frac{1}{s_{k}}\sum_{i}\sum_{j}Z_{ki}Z_{kj}~{}(\mathbf{x}_{i}-\mathbf{x}_{j})^{\intercal}(\mathbf{x}_{i}-\mathbf{x}_{j}).$ (19)
@@
The second energy term $E(Z)$ in Equation 17, modeling the prior distribution over possible assignments cannot exactly be embedded in the Ising model. We therefore use the quadratic penalty method as shows in Equation equation 9 to approximate the energy of an indicator function. Importantly, this does not influence the energy of feasible solutions, as $\lambda||\mathbf{G}\mathbf{x}-\mathbf{d}||_{2}^{2}$ evaluates to zero for all solutions that fulfill the clustering constraints.
4.3 Boltzmann Reparametrization
@@
While ideally the measurements on AQC are direct samples from the Boltzmann distribution (D-Wave Systems Inc., 2017; Dixit et al., 2021), it requires solving a range of challenges that prohibit a direct use in our scenario (Pochart et al., 2021): Mapping between the cost function and the physical system implemented on the AQC requires Hamiltonian scaling (Pochart et al., 2021). Estimating the scaling factor is nontrivial (Nelson et al., 2022) and prohibitive for using samples directly in many cases. Additionally, hardware limitations including imperfection of the processor and the spin-bath polarization effect (Pochart et al., 2021) prevent the AQC to sample the Boltzmann distribution exactly. Finally, as the energy term $E(Z)$ can only be implemented using the penalty method, the sampling density is influenced by solutions not fulfilling the constraints.
We compensate for these limitations by evaluating the energy of all measured feasible solutions $Z^{\prime}$ and recompute $p(Z|\mathbf{X})$ by evaluating the partitioning function over these, which only requires sampling solutions at a sufficiently high temperature.
4.4 Maximum Pointsets
@@
Having the set of the most likely clustering solutions $Z$ available together with their posterior probability $P(Z|\mathbf{X})$ allows to find an assignment $Z^{*}$ of a subset of points that solves the clustering problem with increased probability $P(Z^{*}|\mathbf{X})$. Such a set can be found by using Algorithm 1, which implements a greedy approach that disregards points that disagree between different clustering solutions.
@@
Algorithm 1 MaxsetSearch
In the case of a sufficiently well sampled Boltzmann distribution, the resulting maxset assignment $Z^{*}$ with the corresponding probabilistic estimate $P(Z^{*}|\mathbf{X})$ is still calibrated.
4.5 Inference Parameter Optimization
Using the quadratic penalty method to implement the constraints requires finding suitable Lagrangian multipliers $\lambda$. Even though a very high multiplier theoretically guarantees finding a feasible solution, it also deteriorates the conditioning of the optimization problem. Therefore, a suitable Lagrangian multiplier lifts the cost of any constraint violation above all relevant solutions of the clustering problem, while keeping them low enough to avoid scaling the total energy of the problem up considerably. To estimate the multipliers, we follow an iterative procedure as also proposed in (Zaech et al., 2022).
In an initial step, balanced k-means (Malinen & Fränti, 2014) is used to find a feasible clustering solution. This solution is used to estimate Lagrangian multipliers that avoid any first-order violation of the constraints. In subsequent iterative optimization steps, the problem is solved in simulation, to find multipliers that result in a well conditioned problem.
Such optimization procedure is crucial due to the low fidelity of the current generation of AQCs, which requires careful engineering of the problem energy. Therefore, we expect this procedure to become of reduced importance with the progress of quantum computing.
@@ 5 Experiments and Results
We perform experiments on synthetic data as well as real data to verify the efficacy of our method in finding the set of high-probability solutions and in estimating calibrated confidence scores. The experimental scenarios are solved with QA, Simulated Annealing (SIM), and exact exhaustive search using the presented energy formulation and with k-means as a baseline method, which all optimize for the same cost objective. This further allows us to understand the limitations and required work when deploying the approach to real quantum computers.
5.1 Solver Methods
@@
Due to the strong compute-time limitations on an AQC, all Lagrangian optimization steps are performed with SIM, before measuring the final results on the AQC.
@@
K-Means clustering with a balanced cluster constraint (Malinen & Fränti, 2014) forms the baseline for our approach. We run the algorithm until convergence for a maximum of 1000 iterations. While this solution does not provide a probabilistic estimate, it is useful to assess the relative clustering performance.
5.2 Dataset and Metrics
Data for the quantitative evaluation of our method is synthetically generated. For each clustering problem a total of $I$ points are sampled from a separate normal distribution for each of $K$ clusters. The cluster centers are randomly drawn, such that the distance between each pair of clusters lies within a predefined range $[d_{\text{min}},d_{\text{max}}]$. For each experiment a total of $L$ clustering tasks is generated. This allows us to evaluate the calibration metrics over a large value range. For all experiments that directly compare methods, identical clustering tasks are used.
Further qualitative examples are provided for the IRIS dataset (Fisher, 1936), which contains 50 samples of 4 features in 3 classes. We randomly subsample the points and dimensions to generate the parameters required for our experiments.
Clustering metrics are computed using the available ground-truth clusters. We evaluate 4 standard metrics: The accuracy, which measures the ratio of clustering solutions that are identical to the ground-truth. Completeness (Rosenberg & Hirschberg, 2007) measures the ratio of points from a single cluster being grouped together. The adjusted Rand score (Hubert & Arabie, 1985) compares all pairs of points in the ground truth and prediction and the Fowlkes-Mallows index (Fowlkes & Mallows, 1983) combines precision and recall into a single score.

### v2

@@ I Maximum a-posteriori (MAP):
We now focus on obtaining a single point-estimate as a solution to problem I. To solve the clustering problem using AQC, a QUBO formulation of the K-means energy is required. We use a variation of the one-hot encoding approach [27] to formulate the QUBO of k-means, with cost terms similar to Arthur and Date [4]. It uses a matrix ${Z}\in\{0,1\}^{K\times I}$ to encode the cluster assignment of $I$ samples to $K$ clusters. Each row corresponds to one of $K$ clusters and each column to one of $I$ samples. An entry ${Z}_{ki}=1$ indicates that the sample $x_{i}$ belongs to cluster $c_{k}$. As each sample needs to be assigned to a single cluster, the sum of each column of $Z$ needs to satisfy the constraint $\sum_{k}{Z}_{ki}=1~{}\forall i$. The implementation of constrained clustering, where each cluster has a fixed size $s_{k}$ furthermore requires the row constraints $\sum_{i}{Z}_{ki}=s_{k}~{}\forall k$ on $Z$. The total energy of a solution $E(Z,X)$ can be separated into terms $e(i,j,k)$ modeling the energy of assigning the pair of samples $x_{i}$ and $x_{j}$ to the same cluster $c_{k}$.
$E(Z,X)=\sum_{k}\sum_{i}\sum_{j}Z_{ki}Z_{kj}e(i,j,k)+E(Z)$ (8)
@@
where $E(Z)$ models the row and column constraints as an indicator function. By vectorizing $Z$ in row-major order as ${z}=\text{vec}(Z)$, the energy can be rewritten in matrix form as
$E(Z,X)=E(X|Z)+E(Z)={z}^{\intercal}\,{Q}\,{z}+E(Z),$ (9)
@@
where ${Q}$ is a block diagonal matrix with blocks ${Q}_{0},...,{Q}_{K}$ and ${G}{z}={d}$ corresponds to the matrix formulation of the constraints. Each block ${Q}_{k}$ of ${Q}$ is a square matrix that contains the energy $e(i,j,k)$ at ${Q}_{k,ij}$.
To be solved on the AQC, the potential energy needs to be formulated in a QUBO, which cannot exactly implement $E(Z)$ to model the constraints and is circumvented by using Lagrangian multipliers, leading to the MAP estimate
$\hat{{z}}=\operatorname*{arg\,min}_{z}{z}^{\intercal}{Q^{\prime}}{z}+{b^{\prime}}^{\intercal}{z}$ (10)
@@
with constraints ${Q^{\prime}}={Q}+\lambda{G}^{\intercal}{G}$ and ${b^{\prime}}=-2\lambda{G}^{\intercal}{b}$. To avoid a mixed discrete and continuous optimization problem, a quadratic penalty reformulation $\lambda||{G}{z}-{d}||_{2}^{2}$ is chosen. In contrast to a linear penalty approach, where the multiplier $\lambda$ needs to be optimized, our selection only requires a sufficiently high $\lambda$, as all constraint violations result in an increased penalty term. With such selection, the penalty term evaluates to zero if all constraints are fulfilled, and thus, the minimizer $\hat{{z}}$ of the modified optimization problem is a minimizer of the original optimization problem. Besides modeling balanced clustering, further constraints can be introduced using Lagrangian multipliers. This allows to represent a wide range of tasks as QUBO clustering problems solvable with AQC in the same way.
4.1 Probabilistic Quantum Clustering
The MAP estimate was obtained as the lowest energy solution measured during sampling. We now consider approximating the posterior distribution to quantify the confidence of these solutions matching the actual ground truth. Instead of relying on a sequential Markov Chain Monte Carlo (MCMC) method as done in a plethora of classical algorithms [29], we will approximate the posterior distribution by recomputing the Boltzmann distribution from the set of likely AQC solutions. As the likely solutions are already sampled during quantum annealing, given that the objective follows the EBM, the posterior distribution can be estimated without any large additional computational overhead. Note, this is unlike MCMC, which incurs significant computational load.
@@ 4.1 Probabilistic Quantum Clustering
Our clustering formulation employs a mixture of Gaussians to explain the observations and to define the energy function $E(X|Z)$. Each sample in a cluster $c_{k}$ is modeled as a sample drawn from a Gaussian distribution $\mathcal{N}({\mu}_{k},{I})$ with mean ${\mu}_{k}$ and identity covariance similar to k-means. As shown in Section 4, we are interested in the posterior distribution over possible assignments $Z$:
$p(Z|{X})=\frac{p({X}|{Z})p({Z})}{\sum_{{Z}^{\prime}}p({X}|{Z^{\prime}})}=\frac{p({X}|{Z})p({Z})}{A}.$ (11)
@@
While this approach provides a probabilistic estimate by jointly modeling information about the possible cluster configurations and the distribution of data points, it is often intractable to evaluate due to the partition function ${\sum_{{Z}^{\prime}}p({X}|{Z^{\prime}})}$, the sum over all possible solutions $Z^{\prime}$, which incurs exponential cost in the number of samples. To overcome this, we utilize an AQC that samples directly from the corresponding Boltzmann distribution, which we parameterize according to the probabilistic clustering problem.
Data Model.
@@
Determining the potential energy and thus, cost function of the QUBO is a design choice of the algorithm. For our probabilistic approach, a well-defined data distribution, that forms the basis of the cost function is required. While many tasks approached in quantum computer vision, such as tracking [77] or synchronization [6, 12], costs are based on learned metrics or heuristics, they can also be trained to reflect the properties required in our approach.
@@
We therefore, follow the mixture of Gaussian model, where each cluster generates samples from a normal distribution. With the independence of observations and clusters, given the distribution parameters, the likelihood of the joint observations for an assignment $Z$ is given as the product of the individual likelihoods
$f({X}|{Z})=\prod_{k=1}^{K}f({X}|{Z_{k}})=\prod_{k=1}^{K}\prod_{i\in Z_{k}}f({x}_{i}|{Z_{k}})$ (12)
@@
where the likelihood $f({x}_{i}|{Z_{k}})$ corresponds to a Gaussian distribution that follows $\mathcal{N}({\mu}_{k},{I})$ and $d$ is the dimensionality of the space. This result can be used to formulate the energy-based model with the energy function
$E(X|Z)=\sum_{k=1}^{K}\sum_{i\in Z_{k}}\frac{1}{2}({{x}_{i}}-{{\mu}_{k}})^{\intercal}{I}({{x}_{i}}-{{\mu}_{k}}).$ (13)
@@
$p(Z|{X})=\frac{\mbox{exp}\left[-(E({X}|Z)+E(Z))\right]}{\sum_{{Z}^{\prime}}\mbox{exp}\left[-E({X}|Z^{\prime})\right]}.$ (14)
@@
In this formulation, the energy of the assignment $E(Z)$ corresponds to the prior $p({Z})$, which models feasible and infeasible solutions by an indicator function. In the case of balanced clustering, this corresponds to allowing assignments that have each point assigned to exactly one cluster and a cluster size according to $s_{k}$. When the AQC is used as an optimizer, finding the the lowest energy solution corresponds to the MAP as shown in Equation 7
@@
As the AQC qubit system is an Ising model, it requires formulating $E(X|Z)$ and $E(Z)$ quadratic in the optimization variables $Z$, enabling the joint discovery of assignments and cluster means. This is achieved by using the maximum likelihood estimator of the mean $\mu_{k}=\frac{1}{s_{k}}\sum_{i}Z_{ki}x_{i}$, resulting in a quadratic energy formulation that fits the Ising model in Equation 10
$E_{k}({X}|Z)=\frac{1}{s_{k}}\sum_{i}\sum_{j}Z_{ki}Z_{kj}~{}({x}_{i}-{x}_{j})^{\intercal}({x}_{i}-{x}_{j}).$ (15)
@@
The second energy term $E(Z)$ in Equation 14, modeling the prior distribution over possible assignments cannot exactly be embedded in the Ising model and is approximated using Lagrange multipliers. Importantly, this does not influence the energy of feasible solutions relevant in the posterior distribution, as the penalty evaluates to zero for these.
Posterior recomputation.
@@
While ideally the measurements on AQC are direct samples from the Boltzmann distribution [25, 33], it requires solving a range of challenges that prohibit a direct use in our scenario [68]: Mapping between the cost function and the physical system implemented on the AQC requires Hamiltonian scaling [68]. Estimating the scaling factor is nontrivial [63] and prohibitive for using samples directly in many cases. Additionally, hardware limitations including imperfection of the processor and the spin-bath polarization effect [68] prevent the AQC to sample the Boltzmann distribution exactly. Finally, as the energy term $E(Z)$ can only be implemented using the penalty method, the sampling density is influenced by solutions not fulfilling the constraints. We compensate for these limitations by evaluating the energy of all measured feasible solutions $Z^{\prime}$ and recompute $p(Z|{X})$ by evaluating the analytic partitioning function over these, which only requires sampling solutions at a sufficiently high temperature.
Coresets.
@@
Having the set of the most likely clustering solutions $Z$ available together with their posterior probability $P(Z|X)$ allows to find an assignment $Z^{*}$ of a subset of points that solves the clustering problem with increased probability $P(Z^{*}|X)$. Such a set can be found by using Algorithm 1, which implements a greedy approach that disregards points that disagree between different clustering solutions. With a sufficiently well-sampled Boltzmann distribution, the resulting coreset assignment $Z^{*}$ with the corresponding probabilistic estimate $P(Z^{*}|X)$ is still calibrated.
@@
Algorithm 1 CoresetSearch
Lagrange multiplier optimization.
Using the quadratic penalty method to implement the constraints requires finding suitable Lagrangian multipliers $\lambda$. Even though a very high multiplier theoretically guarantees finding a feasible solution, it also deteriorates the conditioning of the optimization problem. Therefore, a suitable Lagrangian multiplier lifts the cost of any constraint violation above all relevant solutions of the clustering problem, while keeping them low enough to avoid scaling the total energy of the problem up considerably. To estimate the multipliers, we follow an iterative procedure as also proposed in [77].
In an initial step, balanced classical k-means [56] is used to find a feasible clustering solution. This solution is used to estimate Lagrangian multipliers that avoid any first-order violation of the constraints. In subsequent iterative optimization steps, the problem is solved in simulation, to find multipliers that result in a well-conditioned problem.
Such optimization procedure is crucial due to the low fidelity of the current generation of QAs, which requires careful engineering of the problem energy. Therefore, we expect this procedure to become of reduced importance with the progress of quantum computing.
@@ 5 Experiments and Results
We perform experiments on synthetic as well as real data to verify the efficacy of our method in finding the set of high-probability solutions and in estimating calibrated confidence scores. The experimental scenarios are solved with QA, Simulated Annealing (SIM), and exact exhaustive search using the presented energy formulation and with k-means as a baseline method. This further allows us to understand the limitations and required work when deploying the approach to real quantum computers.
Figure 2: Evaluation of the calibration and distribution for QA, SIM and exhaustive search in tasks with 2 to 4 clusters, each with 5 points. All results are generated with 1,000 problems in each scenario and 5,000 measurements for each clustering problem.
Implementation details.
@@
K-Means clustering with a balanced cluster constraint [56] forms the baseline for our approach. We run the algorithm until convergence for a maximum of 1000 iterations. While this solution does not provide a probabilistic estimate, it is useful to assess the relative clustering performance.